# IQA-Eval: Automatic Evaluation of Human-Model Interactive Question Answering

### v1

@@ 5.1 Persona Definitions
As discussed in Section 3.3, we assign personas to LEAs to simulate different groups of humans for diverse human alignments. We investigate assigning the following personas to LEAs, which are defined based on the crowdworker survey results in Lee et al. 24.
@@
Based on the survey results on the distributions of personas of workers 24, for each persona $P$, we split the crowdworkers into two groups: “persons with persona $P$”, and “normal persons without specific persona $P$”. For the first group, we initialize the role prompt in Section 3.3 and use it for the LEAs. For the second group, we utilize default prompting for LEAs. The LEA model in this section is ChatGPT(GPT-3.5-turbo-1106).
@@ 5.2 Experimental Results
Table 3: IQA-Eval evaluation results of IQA models (TDA: TextDavinci; TB: TextBabbage; DA: Davinci). LEAs are assigned specific personas when representing specific groups of workers.
@@
Assigning “Expert” persona leads to a significant decrease in the number of queries because this “Expert” already possesses relevant knowledge and only needs key explanations. The “Clarity-Seeker” requires the highest number of interaction turns among all personas, since it requires additional clarity on questions and entities. Notably, when interacting with TextBabbage and Davinci, the “Clarity-Seeker” achieves the highest accuracy due to a clear understanding obtained from the details of the questions. The results for “Critic-Thinker” and “Adapability-Seeker” in Tables 4 and 4, which rarely improve upon the standard persona, suggest these personas are less common within the overall distributions. It matches our original expectations. In Table 4, the large variation of “# Queries” across different personas indicates that it significantly influences the interaction strategies of LEAs. Furthermore, the accuracy results are around the same level after adding the persona. It demonstrates that the overall performance doesn’t drop with the assignment of personas.
The results in the two tables show that: after assigning persona to LEAs to more accurately represent the specific group of workers, general correlation scores substantially increase. It indicates that assigning specific personas enables LEAs to operate in a more detailed and precise manner, and enhances LEAs to perform IQA-Eval in a more fine-grained and human-aligned way.
@@ 6.1 Datasets
To evaluate the robustness and generalizability of our evaluation framework, we conduct benchmarking across different models on two distinct question answering datasets, each offering unique challenges and complexities requiring advanced reasoning. AmbigQA 35 is a collection of 14,042 annotated questions sourced from the NQ-OPEN benchmarks 22, an open-domain QA dataset. It focuses on questions with inherent ambiguities, reflecting the complexity encountered in real-world queries. These ambiguities often involve diverse aspects such as events, entity references, and answer types, resulting in multiple plausible answers for each question. HotpotQA 53 comprises 113,000 question-answer pairs sourced from Wikipedia, which require multi-hop reasoning spanning multiple documents. It contains a rich array of intricate questions that demand the synthesis of information from various texts to determine accurate answers. In this benchmark, we select 500 questions from each dataset to form a dataset containing 1,000 complex multi-hop and ambiguous questions.
@@ 6.2 LLMs to Benchmark
6.3 Benchmarking Results s
The IQA evaluation benchmarks are presented in Table 5. We divide IQA Models into two categories: weak IQA Models (TextDavinci, TextBabbage, and Davinci) and strong IQA Models (GPT3.5, GPT4, Claude, Llama2, and Zephyr). Weak IQA Models can assist the LEA with answering HotpotQA questions, but due to their knowledge limitations, they cannot help much with AmbigQA questions. Zephyr achieves the lowest performance compared to other strong IQA Models. On the HotpotQA dataset, Zephyr’s accuracy performance is only comparable to the strongest one among weak IQA Models, TextDavinci. Most “Helpfulness” and “Fluency” scores are high (exceeding 3 out of 5), especially for strong IQA Models like GPT4. For the “# of queries”, it is uncommon for interactions to extend beyond two turns. As on HotpotQA, most interactions conclude at the beginning of the second turn, as IQA models have effectively guided users to reach the answers. For AmbigQA, some conversations last longer whereas LEA spends additional turns on clarifying ambiguous entities before approaching the final answer.
@@ 6.4 Analysis
On the more challenging AmbigQA, the stronger model, GPT4, typically requires only one turn to assist LEA in solving questions with high accuracy. In contrast, less powerful models like Llama2 and GPT3.5 need more turns to clarify ambiguous entities, and their accuracy rates are lower. A similar trend is observed on the HotpotQA.
@@
In interaction evaluations, accuracy on final results is not the only metric to show IQA Models’ performance. The quality of intermediate responses is always a significant aspect. On both “helpfulness” and “fluency” metrics, Claude is always the best IQA Model on HotpotQA questions, while on AmbigQA, Llama2 and GPT3.5 outperform GPT4. IQA Model rankings on these two aspects differ from those in Chatbot Arena (non-interaction).
@@
The accuracies in both Table 5 and Table 6 are highly correlated. The Pearson correlations of the accuracy on both datasets are 0.77 and 0.87, respectively. A strong IQA model, such as GPT4, can lead the LEA to finish tasks and largely improve the LEA’s performance on those tasks. Weak assistants may drag down the LEA’s performance, such as the performance of the LEA on both datasets decreases after interacting with Zephyr.
@@ Appendix A Limitations
In conducting this study, certain limitations have influenced our scope and findings. Our methodology was applied exclusively to multi-choice question-answering tasks due to constraints imposed by the datasets used. Moreover, we do not investigate how allowing prompt editing could affect the results. This specialization, while focused, limits the generalizability of our findings across the wider array of question answering formats that exist in both academic research and practical applications.
We advocate for subsequent research efforts to extend the application of our proposed evaluation strategies to a more diverse set of question answering tasks, beyond the multi-choice format. Furthermore, there is a significant opportunity to test these methods with a broader spectrum of LLMs, including those at the cutting edge of the field. Such expansions would not only validate the versatility and robustness of our approaches but also potentially uncover additional insights into the nuances of LLM interaction and performance in varied contexts.
Appendix B Prompts
Both B.1 and B.2 are adapted from Lee et al. [2023]. All the following prompts are from the view of LEA in interactions.
B.1 Interaction Evaluation Prompt
@@
B.2 Interaction Generation Prompt
@@
B.2.1 QA Question and choices format
@@ B.3 Persona Prompts
Table 7: Persona Interaction and Evaluation Descriptions
Appendix C Self-Enhancement Bias
We designed a methodology to further mitigate self-enhancement bias. In this experiment, multiple LEA models evaluate the performance of IQA models during each interaction (“Multi-perspective”). In other words, we introduced third-party evaluations, where various LEA models assess the IQA models’ performance instead of relying solely on the LEA model itself involved in the interaction. After evaluation, we use the average score from all LEA models as the final score. The results of IQA-Eval-Multi-perspective look as in Table 8. The correlations between IQA-Eval-Multi-perspective and human evaluations are in Table 9.
Helpfulness Fluency # Queries Accuracy
LEA models TDA TB DA TDA TB DA TDA TB DA TDA TB DA
Human 4.60 3.84 3.52 4.35 3.84 3.22 1.78 2.57 2.66 0.69 0.52 0.48
LEA models Helpfulness Fluency Overall
GPT4 0.702 0.601 0.624
Claude 0.663 0.613 0.602
GPT3.5 0.641 0.552 0.533
@@ Appendix D Additional Experiments on Natural Questions
We also benchmark IQA models in another dataset called Natural Questions Kwiatkowski et al. [2019b]. This dataset comprises authentic questions posed by users about Wikipedia articles, demanding true multi-turn dialogues for resolution, akin to the setup in multi-turn conversational QA dataset QuAC Choi et al. [2018]. The experiment results are as in Tables 12 and 12. All numbers of queries in the two tables are around 3, and each response to a query from IQA models contains an average of 2 sentences.
Given the number of sentences in each IQA model’s response in Table 12, the non-interactive outputs are roughly equivalent to about two interaction turns, less than three turns in interactive outputs. Thus, The interaction process of IQA-EVAL involves not only reasoning processes but also simulating genuine interactive multi-turn conversations. This suggests that the performances shown in tables 12 and 12 above are driven more by multi-turn interactions than by reasoning processes. Furthermore, these interactions lead to enhanced accuracy, as demonstrated by the superior results in the first two tables compared to those in the last table (non-interactive).
@@ Appendix E Sensitive to Persona Distributions
We expect that IQA-EVAL is sensitive to incorrect persona assignments. We conduct two new experiments to study the effects of changing persona assignments. When the persona distribution is incorrect (such as 20% Expert in Table 13), the performance of IQA-EVAL shows a lower correlation with human evaluations.
@@ Appendix F Accurate QA models are preferred by humans in IQA-EVAL
We also conducted experiments (1) using LEA to evaluate interactions between LEAs and IQA models (interactive) and (2) using LEA to evaluate direct answers generated by IQA models (non-interactive). Our experiments in Tables 15 and 15 show that LEA models prefer accurate models, which aligns well with the conclusion from human annotations.

### v2

@@ 5.1 Persona Definitions
As discussed in Section 3.3, we assign personas to LEAs to simulate different groups of humans for diverse human alignments. We investigate assigning the following personas to LEAs, which are defined based on the crowdworker survey results in Lee et al. 29.
@@
Based on the survey results on the distributions of personas of workers 29, for each persona $P$, we split the crowdworkers into two groups: “persons with persona $P$”, and “normal persons without specific persona $P$”. For the first group, we initialize the role prompt in Section 3.3 and use it for the LEAs. For the second group, we utilize default prompting for LEAs. The LEA model in this section is ChatGPT(GPT-3.5-turbo-1106).
@@ 5.2 Experimental Results
Table 3: IQA-Eval evaluation results of IQA models (TDA: TextDavinci; TB: TextBabbage; DA: Davinci). LEAs, based on GPT3.5, are assigned specific personas when representing specific groups of workers.
@@
The “Expert” persona decrease LEA query counts as “Expert” already possesses relevant knowledge and only needs key explanations. The “Clarity-Seeker” requires the most interaction turns among all personas for comprehension, but achieves the highest accuracy with TextBabbage and Davinci through detailed understanding of questions.
“Critic-Thinker” and “Adapability-Seeker” in Tables 4 and 4 rarely surpass upon the standard persona’s human-preference aligment. We hypothesize that these personas are less reflected within the overall distributions of human preferences. In Table 4, the varying “# Queries” across personas reveals their significant influence on LEA interaction strategies. Accuracy remains consistent after adding personas, showing no performance degradation.
Together, these results indicate that assigning specific personas steer LEAs to perform IQA-Eval in a more fine-grained and human-aligned way.
It is worth nothing that our analysis of persona reassignment shows IQA-Eval is sensitive to incorrect assignments (see Appendix E). Moreover, further analyses about bias evaluation, as well as measuring complementary metrics like offensiveness, are in Appendix G.
@@ 6.1 Datasets
To evaluate the robustness and generalizability of our evaluation framework, we conduct benchmarking across different models on two distinct question answering datasets, each offering unique challenges and complexities requiring advanced reasoning. AmbigQA 42 is a collection of 14,042 annotated questions sourced from the NQ-OPEN benchmarks 27, an open-domain QA dataset. It focuses on questions with inherent ambiguities, reflecting the complexity encountered in real-world queries. These ambiguities often involve diverse aspects such as events, entity references, and answer types, resulting in multiple plausible answers for each question. HotpotQA 62 comprises 113,000 question-answer pairs sourced from Wikipedia, which require multi-hop reasoning spanning multiple documents. It contains a rich array of intricate questions that demand the synthesis of information from various texts to determine accurate answers. In this benchmark, we select 500 questions from each dataset to form a dataset containing 1,000 complex multi-hop and ambiguous questions.
@@ 6.2 LLMs to Benchmark
6.3 Benchmarking Results
The IQA evaluation benchmarks are presented in Table 5. We divide IQA Models into two categories: weak IQA Models (TextDavinci, TextBabbage, and Davinci) and strong IQA Models (GPT3.5, GPT4, Claude, Llama2, and Zephyr). Weak IQA Models can assist the LEA with answering HotpotQA questions, but due to their knowledge limitations, they cannot help much with AmbigQA questions. Zephyr achieves the lowest performance compared to other strong IQA Models. On the HotpotQA dataset, Zephyr’s accuracy performance is only comparable to the strongest one among weak IQA Models, TextDavinci.
Most “Helpfulness” and “Fluency” scores are high (exceeding 3 out of 5), especially for strong IQA Models like GPT4. For the “# of queries”, it is uncommon for interactions to extend beyond two turns. As on HotpotQA, most interactions conclude at the beginning of the second turn, as IQA models have effectively guided users to reach the answers. For AmbigQA, some conversations last longer whereas LEA spends additional turns on clarifying ambiguous entities before approaching the final answer. Additional benchmarking results on the Natural Question dataset are in Appendix D.
6.4 Self-Enhancement Bias
LLMs are shown to demonstrate self-favouring behaviours 44, and no verified or accessible mitigations to this issue exist to the best of our knowledge. This issue is particularly concerning when the LEA models evaluating the IQA models share the same underlying model. In this section, we discuss our two of our attempts to assess the effects of this bias.
Following 66 and 16, we included some empirically useful debiasing instructions as follows:
Table 6: Comparison between new prompts and our prompts used in Table 1. The new prompts are more complex and include effective debiasing instructions.
Helpfulness Fluency Accuracy
LEA models TDA TB DA TDA TB DA TDA TB DA
Human 4.60 3.84 3.52 4.35 3.84 3.22 0.69 0.52 0.48
IQA-EVAL-GPT4 (Our Prompts) 3.67 2.30 2.10 4.77 3.87 3.03 0.87 0.83 0.67
IQA-EVAL-GPT4 (New Prompts) 3.50 2.23 2.10 4.40 4.07 3.53 0.87 0.83 0.67
In Table 5, scores on the row of GPT3.5 is vulnerable to self-enhancement bias. However, with the above debiasing prompt, in Table 7 shows that the results of new prompts are highly similar to the original prompts. Similarly, Table 6 shows that the results of modified and original prompts are differ only lightly.
Table 7: Comparison between new prompts and our prompts used in Table 5 on benchmarking LLMs with IQA-Eval .
LEA models Helpfulness Fluency # Queries Accuracy
IQA-EVAL-GPT3.5 (Our Prompts) 4.72 4.95 1.49 0.63
IQA-EVAL-GPT3.5 (New Prompts) 4.68 4.91 1.35 0.60
We also designed a second methodology to mitigate self-enhancement bias. In this experiment, multiple LEA models evaluate the performance of IQA models during each interaction (“Multi-perspective”). In other words, we introduced third-party evaluations, where various LEA models assess the IQA models’ performance instead of relying solely on the LEA model itself involved in the interaction. After evaluation, we use the average score from all LEA models as the final score. The results of IQA-Eval-Multi-perspective look as in Table 8. The correlations between IQA-Eval-Multi-perspective and human evaluations are in Table 9.
We believe that that the self-preference bias has limited impact on IQA-Eval.
Helpfulness Fluency # Queries Accuracy
LEA models TDA TB DA TDA TB DA TDA TB DA TDA TB DA
Human 4.60 3.84 3.52 4.35 3.84 3.22 1.78 2.57 2.66 0.69 0.52 0.48
LEA models Helpfulness Fluency Overall
IQA-EVAL-GPT4-MP 0.702 0.601 0.624
IQA-EVAL-Claude-MP 0.663 0.613 0.602
IQA-EVAL-GPT3.5-MP 0.641 0.552 0.533
@@ 6.5 Analysis
On the more challenging AmbigQA, the stronger model, GPT4, typically requires only one turn to assist LEA in solving questions with high accuracy. In contrast, less capable models like Llama2 and GPT3.5 need more turns to clarify ambiguous entities and have lower QA accuracies. A similar trend is observed on the HotpotQA.
@@
In interaction evaluations, accuracy on final results is not the only metric to show IQA Models’ performance. The quality of intermediate responses is a significant aspect. On both “helpfulness” and “fluency” metrics, Claude is always the best IQA Model on HotpotQA questions, while on AmbigQA, Llama2 and GPT3.5 outperform GPT4. IQA Model rankings on these two aspects differ from those in Chatbot Arena (non-interaction).
@@
The accuracies in both Table 5 and Table 10 show a consistent trend. The Pearson correlations of the accuracy between the tables are 0.77 and 0.87 on both datasets, respectively. A strong IQA model, such as GPT4, can lead the LEA to finish tasks and largely improve the LEA’s performance on those tasks. Weak assistants may drag down the LEA’s performance, such as the performance of the LEA on both datasets decreases after interacting with Zephyr.
@@ 7 Conclusion
Acknowledgement
We thank the anonymous reviewers for valuable and insightful feedback. This research is supported in part by the National Science Foundation CAREER Grant IIS-2340435, Amazon Research Award and Cisco Research Award. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of the U.S. Government.
@@ Appendix A Limitations
In conducting this study, certain limitations have influenced our scope and findings. First and foremost, LLMs are shown to demonstrate self-favouring behaviours [Panickssery et al., 2024], and no verified or accessible mitigations to this issue exist to the best of our knowledge. We discuss some of our attempts to address this in Section 6.4, but this limitation necessitates future research.
Our methodology was applied exclusively to multi-choice question-answering tasks due to constraints imposed by the datasets used. Moreover, we do not investigate how allowing prompt editing could affect the results. This choice limits the generalizability of our findings across the wider array of question answering formats that exist in both academic research and practical applications.
We advocate for subsequent research efforts to extend the application of our proposed evaluation strategies to a more diverse set of question answering tasks, beyond the multi-choice format.
Furthermore, there is a significant opportunity to test these methods with a broader spectrum of LLMs, including those at the cutting edge of the field. Such expansions would not only validate the versatility and robustness of our approaches but also potentially uncover additional insights into the nuances of LLM interaction and performance in varied contexts.
Appendix B Ethics Statement
In our study, we meticulously crafted each persona to mitigate bias, ensuring they do not adversely impact the IQA-Eval process, as evidenced by the results presented in the table above. However, our personas represent only a limited range. There exists the potential for negative effects from other personas that may be inadequately designed or deliberately biased to achieve specific outcomes. Thus, we oppose irresponsible persona designs that result in biased evaluation results. The principle of persona definition and design should be thoroughly studied in future works. We hope our work will help facilitate future research into better automatic interaction evaluations aligning with crowds.
Appendix C Prompts
Both C.1 and C.2 are adapted from Lee et al. [2023]. All the following prompts are from the view of LEA in interactions.
C.1 Interaction Evaluation Prompt
@@
C.2 Interaction Generation Prompt
@@
C.2.1 QA Question and choices format
@@ Appendix D Additional Experiments on Natural Questions
We benchmark IQA models in another dataset called Natural Questions (Kwiatkowski et al. [2019b]). This dataset comprises authentic questions posed by users about Wikipedia articles, demanding true multi-turn dialogues for resolution, akin to the setup in multi-turn conversational QA dataset QuAC (Choi et al. [2018]). The experiment results are as in Tables 14 and 14. All numbers of queries in the two tables are around 3, and each response to a query from IQA models contains an average of 2 sentences.
Given the number of sentences in each IQA model’s response in Table 14, the non-interactive outputs are roughly equivalent to about two interaction turns, less than three turns in interactive outputs. Thus, The interaction process of IQA-Eval involves not only reasoning processes but also simulating genuine interactive multi-turn conversations. This suggests that the performances shown in tables 14 and 14 above are driven more by multi-turn interactions than by reasoning processes. Furthermore, these interactions lead to enhanced accuracy, as demonstrated by the superior results in the first two tables compared to those in the last table (non-interactive).
@@ Appendix E Sensitive to Persona Distributions
We conduct two new experiments to study the effects of changing persona assignments. Results indicate that IQA-Eval is sensitive to incorrect persona assignments. When the persona distribution is incorrect (such as 20% Expert in Table 15), the performance of IQA-Eval shows a lower correlation with human evaluations.
@@ Appendix F Accurate QA models are preferred by humans in IQA-EVAL
We also conducted experiments (1) using LEA to evaluate interactions between LEAs and IQA models (interactive) and (2) using LEA to evaluate direct answers generated by IQA models (non-interactive). Our experiments in Tables 17 and 17 show that LEA models prefer accurate models, which aligns well with the conclusion from human annotations.
@@
Appendix G Bias Evaluation
We follow the method proposed by Sheng et al. [2021] and conduct a new experiment to evaluate the offensiveness and harmfulness of our personas using the RealToxicityPrompts dataset (Gehman et al. [2020]) on our LEA models. The results are in the table 18. The values in the table above represent the success rates (higher is better) for each bias metric, persona, and LEA model (GPT3.5 and Claude). Scores labeled "None" are consistently lower than those for all personas, indicating that our personas do not increase offensiveness or harmfulness in conversations.
Table 18: Evaluating persona biases on offensiveness and harmful metrics. A high score indicates better results.
Offensiveness Harmful
Persona IQA-EVAL-GPT3.5 IQA-EVAL-Claude IQA-EVAL-GPT3.5 IQA-EVAL-Claude
None 89.5 91.7 62.5 72.5
Expert 95.5 97.3 67.8 75.5
Critical-Thinker 93.3 95.5 65.4 73.7
Adaptability-Seeker 94.5 95.5 62.8 74.6
Clarity-Seeker 95.0 94.0 62.5 73.0